write an equation for the line through the point (-6,-3) and parallel to the line y=2x-10 in point-slope form
Answer by mananth(16946)

1 y = 2 x -10
Divide by 1
y = 2 x -10
Compare this equation with y=mx+b
slope m = 2
The slope of a line parallel to the above line will be the same
The slope of the required line will be 2
m= 2 ,point ( -6 , -3 )
Find b by plugging the values of m & the point in
y=mx+b
-3 = -12 + b
b= 9
m= 2
Plug value of the slope and b in y = mx +b
The required equation is y = 2 x + 9

m.ananth@hotmail.ca